OpenStudy (anonymous):

Find the coordinates of the midpoint of LINE AB if A(-8, 7) and B(-9, 11). quick medal

pooja195 (pooja195):

Same concept \[\huge~\rm~\huge~\rm~\frac{ x_2+x_1 }{ 2},\frac{ y_2+y_1 }{ 2 } \]

pooja195 (pooja195):

Same concept \[\huge~\rm~\huge~\rm~\frac{ -9+-8 }{ 2},\frac{ 11+7 }{ 2 }\]

OpenStudy (anonymous):

\[\frac{ 11+7 }{ 2 } = 9\]

pooja195 (pooja195):

Good now did you get the first one?

OpenStudy (anonymous):

I can't get it

pooja195 (pooja195):

ok -9-8=?

OpenStudy (anonymous):

but it's a negative 8

OpenStudy (alexandervonhumboldt2):

yeah: -9+(-8)=-9-8=?

pooja195 (pooja195):

I know but we are still adding -9+-8 =? bascially says -9-8=?

OpenStudy (anonymous):

\[\frac{ -9-8 }{ 2 } = 8.5\]

OpenStudy (alexandervonhumboldt2):

yeah

pooja195 (pooja195):

or -17/2 :)

OpenStudy (anonymous):

(8.5, 9)

OpenStudy (anonymous):

or (-8.5, 9)

pooja195 (pooja195):

Correct (-8.5, 9)

OpenStudy (anonymous):

thanks, i have more questions i need help on, please help?

pooja195 (pooja195):

Yes :)